# Harmonic $1$-forms on real loci of Calabi-Yau manifolds

### v1

@@ 3.2.3 Intersections of a quadric and a quartic CICY1 and CICY2 with $b_{1}(L)=1$
Let $P_{2}=x_{1}^{2}+x_{2}^{2}$ and $P_{4}=x_{4}^{4}$. Let $\epsilon_{0}\approx 0.590$ be the unique real root of $01-x-2x^{3}=$ 0. Then, for $0<\epsilon<\epsilon_{0}$, the variety $X_{\epsilon}$ is smooth.
@@
00 0 $=|-1+\epsilon x_{1}^{2}+\epsilon x_{2}^{2}+x_{4}^{2}+x_{5}^{2}|\geq 1-\epsilon|x_{1}|^{2}-\epsilon|x_{2}|^{2}\geq 1-2\epsilon^{2}$
which is $>0$ because $\epsilon<\epsilon_{0}<\frac{\sqrt{2}}{2}$. Here we used the triangle inequality for the first inequality and Eq. 3.15 for the second inequality. This is a contradiction.
@@
00 0 $=|-1+\epsilon x_{1}^{2}+\epsilon x_{2}^{2}+x_{4}^{2}+x_{5}^{2}|\geq|x_{4}|^{2}-1-2\epsilon^{2}=\frac{1}{\epsilon}-1-2\epsilon^{2}$
which is $>0$ by the assumption that $\epsilon<\epsilon_{0}$ and the definition of $\epsilon_{0}$. This is again a contradiction.
@@
Note that the singular variety $X^{*}$ has a $\text{U}(1)$-symmetry, namely a rotation in the $(x_{4},x_{5})$-plane. If $X^{*}$ was smooth, it would admit a Calabi-Yau metric, the infinitesimal $\text{U}(1)$-action would define a Killing field, and its dual would be a parallel $1$-form due to $\operatorname{Ric}=0$ and the Bochner trick (see [Bochner, 1946] or [Wu, 2017, Theorem 3.1]). Any smoothing of $X^{*}$ must lose this $\text{U}(1)$-symmetry, as a compact Calabi-Yau manifold with holonomy $\text{SU}(m)$ cannot have a continuous symmetry. However, one may speculate that for small values of $\epsilon$, the infinitesimal $\text{U}(1)$-action projected onto $X_{\epsilon}$ gives rise to a $1$-form that is close to being parallel, and there is a nearby harmonic $1$-form is nowhere vanishing. In Section 5 we present numerical evidence suggesting that this is indeed the case.
@@ 4.1 Sampling algorithms
The algorithms in the following subsections require doing numerical integrals over a Kähler manifold $M$ or its real locus $L$. Consider the integral
@@
and let ${\mathbb{E}}[\cdot]$ be the expectation under sampling from $\operatorname{d}\!\mu_{S}$, then
@@
More generally, we can take the sampling distribution $\operatorname{d}\!\mu_{S}$ different from the volume form $\operatorname{d}\!\mu$. To minimize the sampling variance of $\int\operatorname{d}\!\mu\,f$, we want to choose $\operatorname{d}\!\mu_{S}$ to minimize the variance of $f\,\operatorname{d}\!\mu/\operatorname{d}\!\mu_{S}$. Ideally one would take $\operatorname{d}\!\mu_{S}=f\operatorname{d}\!\mu$ in which case one can take $N=1$. This requires knowing $f$ exactly, which we don’t in our problems. However, the function $f$ is bounded and so the main point is to control the ratio of the volume forms $\operatorname{d}\!\mu/\operatorname{d}\!\mu_{S}$. To do this it is good enough to take $\operatorname{d}\!\mu_{S}$ to be the Fubini-Study metric induced from an embedding into projective space.
@@
This can be used to sample from $M^{d}$ which is a complete intersection on ${\mathbb{C}}\mathbb{P}^{N}$. The direct approach is to choose $d$ random sections on ${\mathbb{C}}\mathbb{P}^{N}$ and intersect their zeroes with those of the defining sections. One can instead dualize and choose $N+1-d$ random points on ${\mathbb{C}}\mathbb{P}^{N}$, whose span is a ${\mathbb{C}}\mathbb{P}^{N-d}$, and take the zeroes of the defining sections on this span.
The method can be used for any ambient space which can be embedded by an explicit complete basis of sections. Below we will use it for products of projective spaces. It was generalized to hypersurfaces in toric varieties in [Braun et al., 2008]. The challenge there is to work out the map between the toric coordinates and the sections; methods for this are discussed in [Larfors et al., 2022].
Next we want to sample points on the real locus $L\subset M$. Note that, up to conjugation, the complex conjugation $z\mapsto\bar{z}$ is the only anti-holomorphic involution on $\mathbb{CP}^{N}$ that has fixed points, see [Morozov, 2022, Lemma 11], so there is little loss of generality in taking the real locus rather than the fixed set of an arbitrary anti-holomorphic involution. A sampling algorithm for real algebraic submanifolds using the technique of intersection with a linear space was developed in [Breiding and Marigliano, 2020] (see also their references [Boykov and Kolmogorov, 2003, Legland et al., 2007, Lehmann and Legland, 2012, Li et al., 2003]). The result most relevant for us is [Breiding and Marigliano, 2020, Theorem 6.1], which we now adapt to our conventions.
@@
One can then approximate this integral by sampling $A$ and averaging over the subsets $x\in L\cap{\mathcal{L}}_{A}$. In words, these are the intersections of linear subspaces of $\mathbb{R}\mathbb{P}^{N}$ with $L$, so this prescription also expresses the measure in terms of zeroes of random sections (here defined by the rows of $A$). One can also sample the same measure by taking ${\mathcal{L}}_{A}$ to be the linear span in $\mathbb{R}^{N+1}$ of $N+1-d$ points sampled from $\operatorname{d}\!\mu_{FS}$. This follows because there is a unique $SO(N+1)$-invariant probability measure on the Grassmannian manifold of subspaces ${\mathcal{L}}$.
To integrate over a manifold embedded in a product of $\mathbb{R}\mathbb{P}^{N}$’s, we can iterate this procedure. The result will be an expectation of a sum of points $x\in L\cap L_{A}\cap L_{B}\ldots\cap L_{X}$, where each subspace $L_{A}$, $L_{B}$ and so on is sampled from one of the projective spaces.
@@ 4.2 Approximating the Calabi-Yau metric using bihomogeneous neural networks
As announced in Section 3.1, we will compute approximate Calabi-Yau metrics on several Calabi-Yau varieties. This problem has a rich history: it was first studied in the physics literature in [Headrick and Wiseman, 2005, Douglas et al., 2008, Braun et al., 2008]. In [Donaldson, 2009] a new approach based on embedding the Calabi-Yau variety into a high-dimensional projective space was pioneered, and this idea still underlies many of the modern approaches. In the subsequent work [Headrick and Nassar, 2013] the numerics were improved, and this approach currently provides the best approximate Calabi-Yau metrics on varieties with many discrete symmetries. For varieties without symmetry, the machine learning approaches [Ashmore et al., 2020, Anderson et al., 2021, Douglas et al., 2022, Douglas, 2022, Jejjala et al., 2022, Larfors et al., 2021, Berglund et al., 2022, Gerdes and Krippendorf, 2023] emerged, and appear to provide the best approximate metrics on such varieties. We also note the work in [Doran et al., 2008, Bunch and Donaldson, 2008, Seyyedali, 2009, Anderson et al., 2010] on very closely related numerical problems.
First, to do a numerically verified version of the analysis in Section 3.1 it is important that the approximate metric be differentiable. (We do not do this analysis in this paper, but hope to do so in the future.) Second, our varieties have few discrete symmetries. These two features restricts the methods we can choose from. We decided to use the bihomogeneous neural networks from [Douglas et al., 2022]. An implementation for hypersurfaces can be found at [Qi, 2022], and we expanded this code to work for complete intersection Calabi-Yaus in a single projective space. A detailed description of the approach can be found in [Douglas et al., 2022, Section 2.4], and we briefly review it here.
Consider a Calabi-Yau manifold $Y\subset\mathbb{CP}^{n}$ defined as a complete intersection of polynomials, such as Fermat or CICY1, and denote coordinates on $\mathbb{CP}^{n}$ by $Z^{0},\dots,Z^{n}$. We will now define a feed-forward neural network
@@
where the $W^{(i)}$’s are real general linear transformations, and $\theta$ is the activation function defined component-wise as $z\mapsto z^{2}$. Here, $Z\bar{Z}$ denotes the real and imaginary parts of all bihomogeneous combinations $Z^{i}\bar{Z}^{j}$, which in complex dimension $n$ are $n^{2}$ inputs. Note that in this way, $F$ is not a well-defined function on $\mathbb{CP}^{n}$, but the Kähler form $\frac{i}{2}\partial\bar{\partial}\log(F)$ is well-defined. That is in analogy to the Kähler form of the ordinary Fubini-Study metric, which is $\omega_{\text{FS}}=\frac{i}{2}\partial\bar{\partial}\log\left(\sum_{i}Z^{i}\bar{Z}^{i}\right)$.
@@ 4.3 Numerical forms using a polynomial basis
Since our surface $L$ is an algebraic hypersurface in real projective space, we can use algebraic techniques to represent and solve for the harmonic one-form. This has two advantages. First, the harmonic form is smooth so a polynomial basis should be highly convergent. Second, it is more similar to our existing code for Calabi-Yau metrics and easier to implement.
Thus, we want to find a harmonic form on $L=Z_{\mathbb{R}}(f)\subset\mathbb{R}\mathbb{P}^{4}$, where the metric $g$ on $L$ is the restriction of a given Kähler metric on the complex hypersurface. (In this section we will work out the details for the quintics. The calculations for the CICYs can be done in the same manner on $\mathbb{R}\mathbb{P}^{5}$.)
@@
Here ||$\cdot$|| ⋅ | is the pointwise norm on $L$ using the metric $g$ which can also be written $|\alpha|^{2}=(\alpha\wedge*_{L}\alpha)/\operatorname{vol}$. Given a parameterized ansatz $\omega_{t}\in\Omega^{1}(L)$ we can get an upper bound on this by minimizing $E(t)$. We then want a numerical estimate for $\min_{x}|\omega(x)|^{2}$.
We have a projection map onto $\langle\operatorname{grad}f\rangle^{\perp}$, say $P_{L}:T\mathbb{RP}^{4}\rightarrow T\mathbb{RP}^{4}$, which induces projections on all forms that we denote by the same symbol by abuse of notation. Given a parameterized ansatz of forms $\omega_{t}\in\Omega^{1}(\mathbb{R}\mathbb{P}^{4})$, the equations Eqs. 4.4 and 4.5 for the form $P_{L}(\omega_{t})|_{L}$ become:
$E_{1}$ $=\int_{L}|P_{L}\,\operatorname{d}\!\,P_{L}\omega|^{2}+|P_{L}\,\operatorname{d}\!\,*_{L}P_{L}\omega|^{2},$
11 11 $=\int_{L}|P_{L}\omega|^{2}.$
The integrals can be estimated by sampling on $L$ and the constraint can be imposed by normalizing $\omega$ in each gradient descent step. Thus the remaining work is to choose a basis and implement $P_{L}$, $*_{L}$ and the norm ||$\cdot$|| ⋅ |.
@@
The $e^{[ij]}$ are an overcomplete basis, i.e. they span $T^{*}_{x}\mathbb{R}\mathbb{P}^{N}$ for all $x\in\mathbb{R}\mathbb{P}^{N}$. To see this, note that at $x_{\text{base}}:=[1:0:\ldots:0]$ the $e^{[1i]}$ with $i>1$ form a basis. The $SO(N+1)$-action then maps the overcomplete basis from $x_{\text{base}}$ to any other point in $\mathbb{R}\mathbb{P}^{N}$. Then $\operatorname{d}\!\,(1/|x|^{2})=-2E/|x|^{2}$ and
@@
It remains to implement $*_{L}$ acting on one-forms and two-forms. The coordinate expression is
$*_{L}\operatorname{d}\!x^{k}=\sqrt{g}\epsilon^{klm}g_{lp}g_{mq}\operatorname{d}\!x^{p}\wedge\operatorname{d}\!x^{q}.$
If we are only using this to write the norm and $*\operatorname{d}\!*\omega$, a bit simpler is to write this as the divergence

### v2

@@ 3.2.3 Intersections of a quadric and a quartic CICY1 and CICY2 with $b_{1}(L)=1$
Let $P_{2}=x_{1}^{2}+x_{2}^{2}$ and $P_{4}=x_{4}^{4}$. Let $\epsilon_{0}\approx 0.590$ be the unique real root of $1-x-2x^{3}=0$. Then, for $0<\epsilon<\epsilon_{0}$, the variety $X_{\epsilon}$ is smooth.
@@
$0$ $=|-1+\epsilon x_{1}^{2}+\epsilon x_{2}^{2}+x_{4}^{2}+x_{5}^{2}|\geq 1-\epsilon|x_{1}|^{2}-\epsilon|x_{2}|^{2}\geq 1-2\epsilon^{2}$
which is >0>0 because $\epsilon<\epsilon_{0}<\frac{\sqrt{2}}{2}$. Here we used the triangle inequality for the first inequality and Eq. 3.15 for the second inequality. This is a contradiction.
@@
$0$ $=|-1+\epsilon x_{1}^{2}+\epsilon x_{2}^{2}+x_{4}^{2}+x_{5}^{2}|\geq|x_{4}|^{2}-1-2\epsilon^{2}=\frac{1}{\epsilon}-1-2\epsilon^{2}$
which is >0>0 by the assumption that $\epsilon<\epsilon_{0}$ and the definition of $\epsilon_{0}$. This is again a contradiction.
@@
Note that the singular variety $X^{*}$ has a $\text{U}(1)$-symmetry, namely a rotation in the $(x_{4},x_{5})$-plane. If $X^{*}$ was smooth, it would admit a Calabi-Yau metric, the infinitesimal $\text{U}(1)$-action would define a Killing field, and its dual would be a parallel $1$-form due to $\operatorname{Ric}=0$ and the Bochner trick (see [Bochner, 1946] or [Wu, 2017, Theorem 3.1]). Any smoothing of $X^{*}$ must lose this $\text{U}(1)$-symmetry, as a compact Calabi-Yau manifold with holonomy $\text{SU}(m)$ cannot have a continuous symmetry. However, one may speculate that for small values of $\epsilon$, the infinitesimal $\text{U}(1)$-action projected onto $X_{\epsilon}$ gives rise to a $1$-form that is close to being parallel, and that there is a nearby harmonic nowhere vanishing $1$-form. In Section 5 we present numerical evidence suggesting that this is indeed the case.
@@ 4.1 Sampling algorithms
The algorithms in the following subsections require doing numerical integrals over a Kähler manifold $M$ or its real locus $L$. For $f:M\rightarrow\mathbb{R}$ and a measure $\operatorname{d}\!\mu_{S}$ on $M$ consider the integral
@@
and let $\mathbb{E}$$\cdot$] be the expectation under sampling from $\operatorname{d}\!\mu_{S}$, then
@@
More generally, we can take the sampling distribution $\operatorname{d}\!\mu_{S}$ different from the Kähler volume form $\operatorname{d}\!\mu$. To minimize the sampling variance of $\int\operatorname{d}\!\mu\,f$, we want to choose $\operatorname{d}\!\mu_{S}$ to minimize the variance of $f\,\operatorname{d}\!\mu/\operatorname{d}\!\mu_{S}$. Ideally one would take $\operatorname{d}\!\mu_{S}=f\operatorname{d}\!\mu$ in which case one can take $N=1$. This requires knowing $f$ exactly, which we don’t in our problems. However, the function $f$ is bounded and so the main point is to control the ratio of the volume forms $\operatorname{d}\!\mu/\operatorname{d}\!\mu_{S}$. To do this it is good enough to take $\operatorname{d}\!\mu_{S}$ to be the Fubini-Study metric induced from an embedding into projective space.
@@
This can be used to sample from $M^{d}$ which is a complete intersection in ${\mathbb{C}}\mathbb{P}^{N}$. The direct approach is to choose $d$ random sections on ${\mathbb{C}}\mathbb{P}^{N}$ and intersect their zeroes with those of the defining sections. One can instead dualise and choose N+1dN+1-d random points on ${\mathbb{C}}\mathbb{P}^{N}$, whose span is a ${\mathbb{C}}\mathbb{P}^{N-d}$, and take the zeroes of the defining sections on this span.
The method can be used for any ambient space which can be embedded by an explicit complete basis of sections. It was generalised to hypersurfaces in toric varieties in [Braun et al., 2008]. The challenge there is to work out the map between the toric coordinates and the sections; methods for this are discussed in [Larfors et al., 2022].
Next we want to sample points on the real locus $L\subset M$. Note that, up to conjugation, the complex conjugation $z\mapsto\bar{z}$ is the only anti-holomorphic involution on $\mathbb{CP}^{N}$ for even $N$, see [Morozov, 2022, Lemma 11], so there is little loss of generality in taking the real locus rather than the fixed set of an arbitrary anti-holomorphic involution. A sampling algorithm for real algebraic submanifolds using the technique of intersection with a linear space was developed in [Breiding and Marigliano, 2020] (see also their references [Boykov and Kolmogorov, 2003, Legland et al., 2007, Lehmann and Legland, 2012, Li et al., 2003]). The result most relevant for us is [Breiding and Marigliano, 2020, Theorem 6.1], which we now adapt to our conventions.
@@
One can then approximate this integral by sampling $A$ and averaging over the subsets $x\in L\cap{\mathcal{L}}_{A}$. In words, these are the intersections of linear subspaces of $\mathbb{R}\mathbb{P}^{N}$ with $L$, so this prescription also expresses the measure in terms of zeroes of random sections (here defined by the rows of $A$). One can also sample the same measure by taking ${\mathcal{L}}_{A}$ to be the linear span in $\mathbb{R}^{N+1}$ of N+1dN+1-d points sampled from $\operatorname{d}\!\mu_{FS}$. This follows because there is a unique $SO(N+1)$-invariant probability measure on the Grassmannian manifold of subspaces $\mathcal{L}$}.
To integrate over a manifold embedded in a product of $\mathbb{R}\mathbb{P}^{N}$’s, one could iterate this procedure. The result would be an expectation of a sum of points $x\in L\cap L_{A}\cap L_{B}\ldots\cap L_{X}$, where each subspace $L_{A}$, $L_{B}$, $\dots$ is sampled from one of the projective spaces.
@@ 4.2 Approximating the Calabi-Yau metric using bihomogeneous neural networks
As announced in Section 3.1, we will compute approximate Calabi-Yau metrics on several Calabi-Yau varieties. This problem has a rich history: it was first studied in [Headrick and Wiseman, 2005] and subsequently [Douglas et al., 2008, Braun et al., 2008]. In [Donaldson, 2009] a new approach based on embedding the Calabi-Yau variety into a high-dimensional projective space was pioneered, and this idea still underlies many of the modern approaches. In the subsequent work [Headrick and Nassar, 2013] the numerics were improved, and this approach currently provides the best approximate Calabi-Yau metrics on varieties with many discrete symmetries. For varieties without symmetry, the machine learning approaches [Ashmore et al., 2020, Anderson et al., 2021, Douglas et al., 2022, Douglas, 2022] as well as [Jejjala et al., 2022, Larfors et al., 2021, Berglund et al., 2022, Gerdes and Krippendorf, 2023] emerged, and appear to provide the best approximate metrics on such varieties. We also note [Doran et al., 2008] as well as [Bunch and Donaldson, 2008, Seyyedali, 2009, Anderson et al., 2010] on very closely related numerical problems.
First, to do a numerically verified version of the analysis in Section 3.1 it is important that the approximate metric be differentiable. (We do not do this analysis in this paper, but hope to do so in the future.) Second, our varieties have few discrete symmetries. These two features restrict the methods we can choose from. We decided to use the bihomogeneous neural networks from [Douglas et al., 2022]. An implementation for hypersurfaces can be found at [Qi, 2022], and we expanded this code to work for complete intersection Calabi-Yaus in a single projective space. A detailed description of the approach can be found in [Douglas et al., 2022, Section 2.4], and we briefly review it here.
Consider a Calabi-Yau manifold $Y\subset\mathbb{CP}^{n}$ defined as a complete intersection of polynomials, such as Fermat or CICY1, and denote coordinates on $\mathbb{CP}^{n}$ by $z_{0},\dots,z_{n}$. We will now define a feed-forward neural network
@@
where the $W^{(i)}$’s are real general linear transformations, and $\theta$ is the activation function defined in every component as $z\mapsto z^{2}$. Here, $z\bar{z}$ denotes the real and imaginary parts of all bihomogeneous combinations $z_{i}\bar{z}_{j}$, which in complex dimension $n$ are $n^{2}$ inputs. Note that in this way, $F$ is not a well-defined function on $\mathbb{CP}^{n}$, but the Kähler form $\frac{i}{2}\partial\bar{\partial}\log(F)$ is well-defined. That is in analogy to the Kähler form of the ordinary Fubini-Study metric, which is $\omega_{\text{FS}}=\frac{i}{2}\partial\bar{\partial}\log\left(\sum_{i}z_{i}\bar{z}_{i}\right)$.
@@ 4.3 Numerical forms using a polynomial basis
Since our surface $L$ is an algebraic hypersurface in real projective space, we can use algebraic techniques to represent and solve for the harmonic one-form. This has two advantages. First, the harmonic form is smooth so a polynomial basis should be highly convergent. Second, it is similar to our existing code for Calabi-Yau metrics and easy to implement.
Thus, we want to find a harmonic form on $L=Z_{\mathbb{R}}(f)\subset\mathbb{R}\mathbb{P}^{4}$, where the metric $g$ on $L$ is the restriction of a given Kähler metric on the complex hypersurface. (In this section we will work out the details for quintic hypersurfaces. The calculations for CICYs can be done in the same manner on $\mathbb{R}\mathbb{P}^{5}$.)
@@
Here $|\cdot|$ is the pointwise norm on $L$ using the metric $g$ which can also be written $|\alpha|^{2}=(\alpha\wedge*_{L}\alpha)/\operatorname{vol}$. Given a parameterized ansatz $\omega_{t}\in\Omega^{1}(L)$ we can get an upper bound on this by minimizing $E(t)$.
We have a projection map onto $\langle\operatorname{grad}f\rangle^{\perp}$, say $P_{L}:T\mathbb{RP}^{4}\rightarrow T\mathbb{RP}^{4}$, which induces projections on all differential forms that we denote by the same symbol by abuse of notation. Given a parameterized ansatz of forms $\omega_{t}\in\Omega^{1}(\mathbb{R}\mathbb{P}^{4})$, the equations Eqs. 4.4 and 4.5 for the form $P_{L}(\omega_{t})|_{L}$ become:
$E_{1}$ $=\int_{L}|P_{L}\,\operatorname{d}\!\,P_{L}\omega_{t}|^{2}+|P_{L}\,\operatorname{d}\!\,*_{L}P_{L}\omega_{t}|^{2},$
$1$ $=\int_{L}|P_{L}\omega_{t}|^{2}.$
The integrals can be estimated by sampling on $L$ and the constraint can be imposed by normalizing $\omega$ in each gradient descent step. Thus the remaining work is to choose a basis and implement $P_{L}$, $_{L}$ and the norm $|\cdot|$.
@@
The $e^{[ij]}$ are an overcomplete basis, i.e. they span $T^{*}_{x}\mathbb{R}\mathbb{P}^{N}$ for all $x\in\mathbb{R}\mathbb{P}^{N}$. To see this, note that at $x_{\text{base}}:=[1:0:\ldots:0]$ the $e^{[1i]}$ with $i>1$ form a basis. The $SO(N+1)$-action then maps the overcomplete basis from $x_{\text{base}}$ to any other point in $\mathbb{R}\mathbb{P}^{N}$. Furthermore $\operatorname{d}\!\,(1/|x|^{2})=-2E/|x|^{2}$ and
@@
It remains to implement $_{L}$ acting on one-forms and two-forms. The coordinate expression is
$_{L}\operatorname{d}\!x^{k}=\sqrt{g}\epsilon^{klm}g_{lp}g_{mq}\operatorname{d}\!x^{p}\wedge\operatorname{d}\!x^{q}.$
If we are only using this to write the norm and $\operatorname{d}\!*\omega$, a bit simpler is to write this as the divergence